# CryptoMoE: Privacy-Preserving and Scalable Mixture of Experts Inference via Balanced Expert Routing

### v1

@@ 1 Introduction
Sparsely-gated mixture-of-expert (MoE) models have emerged as a powerful architecture for scaling up large language model (LLM) capacity without proportionally increasing the computation cost. As a result, many state-of-the-art LLM families, including LLaMA-4 [1], DeepSeek-V3 [2], and QWen-3 [3], have adopted MoE as their core architecture.
Driven by the high model capacity, MoE-based LLMs are increasingly adopted in real-world applications, some of which involve sensitive user data, e.g., person re-identification [4] and medical diagnostics [5]. Therefore, data privacy has become a major concern and has propelled the development of privacy-preserving inference frameworks. Hybrid cryptographic approaches combining Homomorphic Encryption (HE) and Secure Multi-Party Computation (MPC) are considered a promising solution. They enable the user and model provider (server) to jointly compute LLM outputs without exposing either the user inputs or the model weights.
However, existing private inference frameworks primarily support dense architectures such as GPT-2 [6] and LLaMA-1 [7], and lack support for MoE-based models. A core challenge lies in how to securely evaluate the dynamic routing mechanism inherent to MoE layers. As illustrated in Figure 1(a), MoE operates by activating a subset of experts for each input token, where each expert is a distinct sub-network. Figure 1(b) further demonstrates that expert activation patterns are highly input-dependent and need to be protected: in 10-th layer of DeepSeekMoE [8], experts #3 and #60 are frequently activated for mathematical tasks across two math datasets [9, 10], but exhibit a uniform distribution across eight textual reasoning datasets [11, 12, 13, 14, 15, 16, 17]. Similar findings have been reported in previous work [18, 19]. This indicates that individual experts often specialize in specific semantic domains. Consequently, revealing expert routing information may leak sensitive details about both the input type and the internal specialization of different experts.
@@
We evaluate CryptoMoE on three representative MoE-based LLMs: DeepSeekMoE-2.8B/16.4B[8], QWenMoE-2.7B/14.3B [20], and OLMoE-1.3B/6.9B [21], across eight zero-shot reasoning tasks. Results show that CryptoMoE retains 99.2% of the original accuracy on average, while achieving a $2.8\sim 3.5\times$ speedup over dense baseline. Moreover, CryptoMoE achieves efficiency comparable to the insecure baseline that fully reveals routing information. We also adapt CipherPrune [22]’s pruning protocol to the MoE setting and construct a strong baseline. Compared to it,CryptoMoE achieves up to $4.3\times$ communication reduction and $2.4\times$ latency reduction. These results establish CryptoMoE as the first framework to enable private, efficient, and accurate inference for MoE-based LLMs.
@@ 2.1 Mixture of Experts Layer
Each expert network $E_{i}$ is a feed-forward network (FFN) implemented by SwiGLU [23]. For input with multiple tokens $\{x_{i}\}_{i=0}^{m-1}$, tokens are routed to different experts based on the gating network. Then each expert network $E_{i}$ processes the tokens distributed to it in parallel.
@@ 2.2 Cryptographic Primitives
Homomorphic Encryption (HE). Following most hybrid HE/MPC schemes [24, 25, 26, 27, 28], CryptoMoE leverages the Brakerski-Fan-Vercauteren (BFV) HE scheme [29] and mainly involves the following element-wise HE operations: ciphertext addition, ciphertext-plaintext multiplication, and ciphertext rotation $\operatorname{Rot}(\operatorname{ct},s)$, which shifts the ciphertext $\operatorname{ct}$ to the left by $s$ positions.
Secure Multi-Party Computation (MPC). We employ a 2-out-of-2 additive Secret Share (SS)-based MPC scheme [30] to keep the input data private throughout inference. We denote two parties by $P_{0}$ and $P_{1}$, where $P_{0}$ is the client and $P_{1}$ is the server. We use [[x]][\![x]\!] to denote an additive share of $x$. We write $x]\!]=([\![x]\!]_{0},[\![x]\!]_{1})$ where $P_{0}$ holds $x]\!]_{0}$ and $P_{1}$ holds $x]\!]_{1}$, such that $x]\!]_{0}+[\![x]\!]_{1}=x$. We write $x]\!]^{B}$ to denote the share of Boolean data. This work builds upon pre-existing MPC protocols whose input and output are additive shares [30, 24, 31]. These protocols are summarized in Table 1. Among them, $\Pi_{\text{matmul}}$ is implemented using HE [24], while the rest are implemented using oblivious transfer (OT) [30, 31].
Threat Model and Security Guarantee. CryptoMoE works in a general private inference scenario that involves two parties, i.e., server $P_{1}$ and client $P_{0}$. The server holds the proprietary NN model, and the client owns private input [26, 28, 27, 32, 24, 31]. CryptoMoE enables the client to obtain the inference results while keeping the server’s model weights and the client’s input private. Consistent with previous works [26, 30, 33, 27, 24, 31], CryptoMoE adopts an honest-but-curious security model in which both parties follow the specification of the protocol but also try to learn more than allowed. CryptoMoE is built upon cryptographic primitives, including BFV and MPC protocols, the security can hence be guaranteed following [29, 34].
@@ 2.3 Related Work
With the proliferation of ChatGPT, significant efforts have been made to enable private Transformer inference, including hybrid HE/MPC frameworks [32, 24, 31, 35, 36, 37, 38], Fully-HE frameworks [39, 40, 41] and Fully-MPC frameworks [42, 43, 44, 45]. However, these works only support dense models like GPT-2 [6]. Recent work, CipherPrune [22], introduces dynamic token pruning for private Transformer inference. Nevertheless, applying its pruning protocol directly to MoE layers not only leaks the number of tokens assigned to each expert but also incurs substantial communication overhead. Thus, how to support private MoE-based model inference is still an open question.
@@ 4.1 Inference-Time Balanced Expert Routing
Challenge 2: Construction of $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$. While step ❶❸ can be implemented by existing protocol $\Pi_{\text{Softmax}},\Pi_{\text{MatMul}}$ and $\Pi_{\text{Top-K}}$ proposed in Bolt [24] and Bumblebee [31], constructing $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ is non-trivial. $\Pi_{\text{dispatch}}$ must securely assign tokens to experts based on routing information and select the $t$ tokens for each expert. Similarly, constructing the $\Pi_{\text{combine}}$ protocol to aggregate expert outputs into token-wise results is complex. Designing both protocols using MPC and HE protocols to preserve privacy without incurring significant overhead remains a challenge.
@@ 4.2 Confidence-Aware Secure Dispatch Protocol
Compute token priority scores. Use $\Pi_{\text{equal}}$ to evaluate whether selected expert indices [[K]][\![K]\!] are equal to $i$, producing boolean mask $M_{i}]\!]\in\{0,1\}^{km}$ that indicate which tokens are useful for this expert. Next, we employ $\Pi_{\text{mux}}$ to combine the routing scores [[W]][\![W]\!] with the masks $M_{i}]\!]$, resulting in token priority scores $S_{i}]\!]$. Useful tokens retain their original scores, while useless (dummy) tokens’ scores are set to zero. To proceed, we must select the top $t$ tokens from the $km$ candidates based on these scores. Prior work CipherPrune [22], introduces a secure pruning protocol capable of achieving this. However, it incurs a high communication cost of $O(kmtd)$, where $d$ denotes the hidden dimension. This overhead arises because each comparison involves a secure swap of the whole token embeddings, making it impractical for MoE inference. To address this limitation, we propose a novel protocol that decouples token scores from embeddings, thereby reducing the communication complexity to $O(km\log(km))$. A detailed comparison with CipherPrune is provided in Appendix C. Our protocol comprises two steps: ❷ Select token indices. We apply $\Pi_{\text{topk}}(k=t)$ to select the $t$ tokens with the highest priority scores. The resulting indices $K_{i}]\!]\in\mathbb{Z}^{t}$ correspond to token positions within the dispatched sequence of length $km$. To map these indices back to the original $m$-token input sequence, we perform an integer division by $k$ to get $K_{i}^{\prime}]\!]=[\![\lfloor\frac{K_{i}}{k}\rfloor]\!]$. ❸ Retrieve tokens. Using the selected indices, we retrieve the corresponding token embeddings. Specifically, we convert the selected indices $K^{\prime}_{i}]\!]\in\mathbb{Z}^{t}$ into a one-hot matrix using $\Pi_{\text{onehot}}([\![K^{\prime}_{i}]\!],m)\in\mathbb{Z}^{t\times m}$, which requires $O(tm)$ calls to $\Pi_{\text{equal}}$. We then perform a $\Pi_{\text{MatMul}}$ with the input token embeddings $x]\!]\in\mathbb{Z}^{m\times d}$ to obtain the $t$ desired token embeddings for expert $i$, i.e., $\mathcal{X}_{i}]\!]\in\mathbb{Z}^{t\times d}$.
@@ 4.3 Efficient Secure Combine Protocol
Complexity Analysis. For a single MoE layer, the proposed $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ introduce additional communication overhead of $O(nkm\log(km)+ntm)$, where the first term stems from the $\Pi_{\text{topk}}$ protocol [46] and the second from $\Pi_{\text{onehot}}$. Experimental results show that our protocol is highly efficient, incurring only an 18% overhead while preserving privacy and leveraging the sparsity of MoE computation. Further implementation details of $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ are provided in Appendix B.
@@ 4.4 Efficient Batch Ciphertext-Plaintext MatMul (Batch MatMul) Protocol
Linear layer evaluation remains a major bottleneck in HE for expert computation. Unlike dense models, MoE layers process $n$ groups of parallel tokens, each of size $t\times d_{1}$ ($t$ tokens, each with an embedding dimension of $d_{1}$). Existing packing schemes like BOLT [24] optimize for dense models by packing along the $t$-dimension, reducing the packed hidden size in a ciphertext and thus minimizing expensive ciphertext rotations. However, in MoE layers, each expert handles only a few tokens. Applying these schemes increases the packed hidden dimension, leading to more rotations and an extremely higher computation cost.
To address this, we introduce an efficient batch ciphertext-plaintext MatMul protocol tailored for MoE computation. The key idea is to pack partial token embeddings from all experts into a single ciphertext. Figure 6 shows a toy example where $n=2$, $t=2$, $d_{1}=4$, with each ciphertext holding 8 elements. In the original packing scheme, each expert’s input matrix of size $t\times d_{1}$ is packed into a ciphertext, requiring 3 rotations per MatMul to accumulate partial sums, resulting in 6 total rotations for two experts. In contrast, our method packs partial embeddings of all tokens into a single ciphertext with shape $(nt\times\frac{d_{1}}{n})$. The weight matrices are adjusted accordingly in plaintext without additional overhead. As shown in Figure 6 (b), this reduces the number of rotations to just 2, thanks to the smaller hidden dimension in each ciphertext. Our batch MatMul protocol reduces HE rotations from $O(nd_{1})$ to $O(d_{1})$. Our method is also compatible with the Baby-Step Giant-Step (BSGS) algorithm [24], which can further reduce the number of rotation operations. The Complexity analysis is provided in Appendix D.
@@ 5.1 Experimental Setup
Implementation. We implement CryptoMoE upon the SecretFlow-SPU framework [47], which is a popular framework for privacy-preserving deep learning. We adopt a secure two-party computation (2PC) setting without a trusted third party. All the experiments are performed on a machine with an Intel Xeon Platinum 8468 CPU (48 cores and 2.1GHz). We consider two network environments: 1) LAN setting with 3Gbps bandwidth and 0.2ms latency; 2) WAN setting with 400Mbps bandwidth and 40ms latency. We simulate network environment via Linux Traffic Control.
Datasets and Models. We consider three popular MoE models: 1) DeepSeekMoE-2.8B/16.4B (6 of 64 experts) [8], 2) QWenMoE-2.7B/14.3B (4 of 60 experts) [20] and 3) OLMoE-1.3B/6.9B (6 of 60 experts) (ICLR’25 Oral) [21]. Since MoE models typically follow a similar design, CryptoMoE can also be applied to other MoE models. All the models are evaluated on eight famous zero-shot common sense reasoning tasks, including SIQA [13], OBQA [17], BoolQ [11], ARC-easy, ARC-challenge [16], HellaSwag [14], PIQA [12], and WinoGrande [15].
Baselines. Since CryptoMoE is the first framework enabling private MoE inference, we compare it with three baselines: 1) Insecure baseline, 2) Dense baseline, 3) CryptoMoE baseline where $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ are constructed by CipherPrune [22]’s pruning protocol.

### v3

@@ 1 Introduction
Sparsely-gated mixture-of-expert (MoE) models have emerged as a powerful architecture for scaling up large language model (LLM) capacity without proportionally increasing the computation cost. As a result, many state-of-the-art LLM families, including LLaMA-4 [LLAMA4], DeepSeek-V3 [liu2024deepseekv3], and QWen-3 [qwen2.5], have adopted MoE as their core architecture.
Driven by the high model capacity, MoE-based LLMs are increasingly adopted in real-world applications, some of which involve sensitive user data, e.g., person re-identification [sarker2024transformer] and medical diagnostics [shamshad2023Transformers]. Therefore, data privacy has become a major concern and has propelled the development of privacy-preserving inference frameworks. Hybrid cryptographic approaches combining Homomorphic Encryption (HE) and Secure Multi-Party Computation (MPC) are considered a promising solution. They enable the user and model provider (server) to jointly compute LLM outputs without exposing either the user inputs or the model weights.
However, existing private inference frameworks primarily support dense architectures such as GPT-2 [radford2019languagegpt2] and LLaMA-1 [touvron2023llama], and lack support for MoE-based models. A core challenge lies in how to securely evaluate the dynamic routing mechanism inherent to MoE layers. As illustrated in Figure 1(a), MoE operates by activating a subset of experts for each input token, where each expert is a distinct sub-network. Figure 1(b) further demonstrates that expert activation patterns are highly input-dependent and need to be protected: in 10-th layer of DeepSeekMoE [dai2024deepseekmoe], experts #3 and #60 are frequently activated for mathematical tasks across two math datasets [gsm8k, amini2019mathqa], but exhibit a uniform distribution across eight textual reasoning datasets [clark2019boolq, bisk2020piqa, sap2019socialiqa, zellers2019hellaswag, sakaguchi2021winogrande, arceasy_challenge, OBQA]. Similar findings have been reported in previous work [yang2025mixture, li2024locmoe]. This indicates that individual experts often specialize in specific semantic domains. Consequently, revealing expert routing information may leak sensitive details about both the input type and the internal specialization of different experts.
@@
We evaluate CryptoMoE on three representative MoE-based LLMs: DeepSeekMoE-2.8B/16.4B[dai2024deepseekmoe], QWenMoE-2.7B/14.3B [bai2023qwen1.5], and OLMoE-1.3B/6.9B [muennighoff2024olmoe], across eight zero-shot reasoning tasks. Results show that CryptoMoE retains 99.2% of the original accuracy on average, while achieving a $2.8\sim 3.5\times$ speedup over dense baseline. Moreover, CryptoMoE achieves efficiency comparable to the insecure baseline that fully reveals routing information. We also adapt CipherPrune [zhang2025cipherprune]’s pruning protocol to the MoE setting and construct a strong baseline. Compared to it,CryptoMoE achieves up to $4.3\times$ communication reduction and $2.4\times$ latency reduction. These results establish CryptoMoE as the first framework to enable private, efficient, and accurate inference for MoE-based LLMs.
@@ 2.1 Mixture of Experts Layer
Each expert network $E_{i}$ is a feed-forward network (FFN) implemented by SwiGLU [shazeer2020glu]. For input with multiple tokens $\{x_{i}\}_{i=0}^{m-1}$, tokens are routed to different experts based on the gating network. Then each expert network $E_{i}$ processes the tokens distributed to it in parallel.
@@ 2.2 Cryptographic Primitives
Homomorphic Encryption (HE). Following most hybrid HE/MPC schemes [pang2023bolt, gilad2016cryptonets, Juvekar_Vaikuntanathan_gazelle_2018, huang2022cheetah, Mishra_Delphi_2020], CryptoMoE leverages the Brakerski-Fan-Vercauteren (BFV) HE scheme [fan2012somewhat] and mainly involves the following element-wise HE operations: ciphertext addition, ciphertext-plaintext multiplication, and ciphertext rotation $\operatorname{Rot}(\operatorname{ct},s)$, which shifts the ciphertext $\operatorname{ct}$ to the left by $s$ positions.
Secure Multi-Party Computation (MPC). We employ a 2-out-of-2 additive Secret Share (SS)-based MPC scheme [rathee2020cryptflow2] to keep the input data private throughout inference. We denote two parties by $P_{0}$ and $P_{1}$, where $P_{0}$ is the client and $P_{1}$ is the server. We use [[x]][\![x]\!] to denote an additive share of $x$. We write $x]\!]=([\![x]\!]_{0},[\![x]\!]_{1})$ where $P_{0}$ holds $x]\!]_{0}$ and $P_{1}$ holds $x]\!]_{1}$, such that $x]\!]_{0}+[\![x]\!]_{1}=x$. We write $x]\!]^{B}$ to denote the share of Boolean data. This work builds upon pre-existing MPC protocols whose input and output are additive shares [rathee2020cryptflow2, pang2023bolt, lu2023bumblebee]. These protocols are summarized in Table 1. Among them, $\Pi_{\text{matmul}}$ is implemented using HE [pang2023bolt], while the rest are implemented using oblivious transfer (OT) [rathee2020cryptflow2, lu2023bumblebee].
Threat Model and Security Guarantee. CryptoMoE works in a general private inference scenario that involves two parties, i.e., server $P_{1}$ and client $P_{0}$. The server holds the proprietary NN model, and the client owns private input [Juvekar_Vaikuntanathan_gazelle_2018, Mishra_Delphi_2020, huang2022cheetah, hao2022iron, pang2023bolt, lu2023bumblebee]. CryptoMoE enables the client to obtain the inference results while keeping the server’s model weights and the client’s input private. Consistent with previous works [Juvekar_Vaikuntanathan_gazelle_2018, rathee2020cryptflow2, rathee2021sirnn, huang2022cheetah, pang2023bolt, lu2023bumblebee], CryptoMoE adopts an honest-but-curious security model in which both parties follow the specification of the protocol but also try to learn more than allowed. CryptoMoE is built upon cryptographic primitives, including BFV and MPC protocols, the security can hence be guaranteed following [fan2012somewhat, goldreich1998secure].
@@ 2.3 Related Work
With the proliferation of ChatGPT, significant efforts have been made to enable private Transformer inference, including hybrid HE/MPC frameworks [hao2022iron, pang2023bolt, lu2023bumblebee, huang2024secbert, luo2024secformer, zeng2024securegpt, xu2024privcirnet], Fully-HE frameworks [cryptoeprint:2024/136NEXUS, park2024powerformer, cryptoeprint:2024/1881THOR] and Fully-MPC frameworks [gupta2023sigma, dong2023puma, akimoto2023privformer, li2022mpcformer]. However, these works only support dense models like GPT-2 [radford2019languagegpt2]. Recent work, CipherPrune [zhang2025cipherprune], introduces dynamic token pruning for private Transformer inference. Nevertheless, applying its pruning protocol directly to MoE layers not only leaks the number of tokens assigned to each expert but also incurs substantial communication overhead. Thus, how to support private MoE-based model inference is still an open question.
@@ 4.1 Inference-Time Balanced Expert Routing
Challenge 2: Construction of $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$. While step ❶❸ can be implemented by existing protocol $\Pi_{\text{Softmax}},\Pi_{\text{MatMul}}$ and $\Pi_{\text{Top-K}}$ proposed in Bolt [pang2023bolt] and Bumblebee [lu2023bumblebee], constructing $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ is non-trivial. $\Pi_{\text{dispatch}}$ must securely assign tokens to experts based on routing information and select the $t$ tokens for each expert. Similarly, constructing the $\Pi_{\text{combine}}$ protocol to aggregate expert outputs into token-wise results is complex. Designing both protocols using MPC and HE protocols to preserve privacy without incurring significant overhead remains a challenge.
@@ 4.2 Confidence-Aware Secure Dispatch Protocol
Compute token priority scores. Use $\Pi_{\text{equal}}$ to evaluate whether selected expert indices [[K]][\![K]\!] are equal to $i$, producing boolean mask $M_{i}]\!]\in\{0,1\}^{km}$ that indicate which tokens are useful for this expert. Next, we employ $\Pi_{\text{mux}}$ to combine the routing scores [[W]][\![W]\!] with the masks $M_{i}]\!]$, resulting in token priority scores $S_{i}]\!]$. Useful tokens retain their original scores, while useless (dummy) tokens’ scores are set to zero. To proceed, we must select the top $t$ tokens from the $km$ candidates based on these scores. Prior work CipherPrune [zhang2025cipherprune], introduces a secure pruning protocol capable of achieving this. However, it incurs a high communication cost of $O(kmtd)$, where $d$ denotes the hidden dimension. This overhead arises because each comparison involves a secure swap of the whole token embeddings, making it impractical for MoE inference. To address this limitation, we propose a novel protocol that decouples token scores from embeddings, thereby reducing the communication complexity to $O(km\log(km))$. A detailed comparison with CipherPrune is provided in Appendix C. Our protocol comprises two steps: ❷ Select token indices. We apply $\Pi_{\text{topk}}(k=t)$ to select the $t$ tokens with the highest priority scores. The resulting indices $K_{i}]\!]\in\mathbb{Z}^{t}$ correspond to token positions within the dispatched sequence of length $km$. To map these indices back to the original $m$-token input sequence, we perform an integer division by $k$ to get $K_{i}^{\prime}]\!]=[\![\lfloor\frac{K_{i}}{k}\rfloor]\!]$. ❸ Retrieve tokens. Using the selected indices, we retrieve the corresponding token embeddings. Specifically, we convert the selected indices $K^{\prime}_{i}]\!]\in\mathbb{Z}^{t}$ into a one-hot matrix using $\Pi_{\text{onehot}}([\![K^{\prime}_{i}]\!],m)\in\mathbb{Z}^{t\times m}$, which requires $O(tm)$ calls to $\Pi_{\text{equal}}$. We then perform a $\Pi_{\text{MatMul}}$ with the input token embeddings $x]\!]\in\mathbb{Z}^{m\times d}$ to obtain the $t$ desired token embeddings for expert $i$, i.e., $\mathcal{X}_{i}]\!]\in\mathbb{Z}^{t\times d}$.
@@ 4.3 Efficient Secure Combine Protocol
Complexity Analysis. For a single MoE layer, the proposed $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ introduce additional communication overhead of $O(nkm\log(km)+ntm)$, where the first term stems from the $\Pi_{\text{topk}}$ protocol [hou2023ciphergpt] and the second from $\Pi_{\text{onehot}}$. Experimental results show that our protocol is highly efficient, incurring only an 18% overhead while preserving privacy and leveraging the sparsity of MoE computation. Further implementation details of $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ are provided in Appendix B.
@@ 4.4 Efficient Batch Ciphertext-Plaintext MatMul (Batch MatMul) Protocol
Linear layer evaluation remains a major bottleneck in HE for expert computation. Unlike dense models, MoE layers process $n$ groups of parallel tokens, each of size $t\times d_{1}$ ($t$ tokens, each with an embedding dimension of $d_{1}$). Existing packing schemes like BOLT [pang2023bolt] optimize for dense models by packing along the $t$-dimension, reducing the packed hidden size in a ciphertext and thus minimizing expensive ciphertext rotations. However, in MoE layers, each expert handles only a few tokens. Applying these schemes increases the packed hidden dimension, leading to more rotations and an extremely higher computation cost.
To address this, we introduce an efficient batch ciphertext-plaintext MatMul protocol tailored for MoE computation. The key idea is to pack partial token embeddings from all experts into a single ciphertext. Figure 6 shows a toy example where $n=2$, $t=2$, $d_{1}=4$, with each ciphertext holding 8 elements. In the original packing scheme, each expert’s input matrix of size $t\times d_{1}$ is packed into a ciphertext, requiring 3 rotations per MatMul to accumulate partial sums, resulting in 6 total rotations for two experts. In contrast, our method packs partial embeddings of all tokens into a single ciphertext with shape $(nt\times\frac{d_{1}}{n})$. The weight matrices are adjusted accordingly in plaintext without additional overhead. As shown in Figure 6 (b), this reduces the number of rotations to just 2, thanks to the smaller hidden dimension in each ciphertext. Our batch MatMul protocol reduces HE rotations from $O(nd_{1})$ to $O(d_{1})$. Our method is also compatible with the Baby-Step Giant-Step (BSGS) algorithm [pang2023bolt], which can further reduce the number of rotation operations. The Complexity analysis is provided in Appendix D.
@@ 5.1 Experimental Setup
Implementation. We implement CryptoMoE upon the SecretFlow-SPU framework [spu], which is a popular framework for privacy-preserving deep learning. We adopt a secure two-party computation (2PC) setting without a trusted third party. All the experiments are performed on a machine with an Intel Xeon Platinum 8468 CPU (48 cores and 2.1GHz). We consider two network environments: 1) LAN setting with 3Gbps bandwidth and 0.2ms latency; 2) WAN setting with 400Mbps bandwidth and 40ms latency. We simulate network environment via Linux Traffic Control.
Datasets and Models. We consider three popular MoE models: 1) DeepSeekMoE-2.8B/16.4B (6 of 64 experts) [dai2024deepseekmoe], 2) QWenMoE-2.7B/14.3B (4 of 60 experts) [bai2023qwen1.5] and 3) OLMoE-1.3B/6.9B (6 of 60 experts) (ICLR’25 Oral) [muennighoff2024olmoe]. Since MoE models typically follow a similar design, CryptoMoE can also be applied to other MoE models. All the models are evaluated on eight famous zero-shot common sense reasoning tasks, including SIQA [sap2019socialiqa], OBQA [OBQA], BoolQ [clark2019boolq], ARC-easy, ARC-challenge [arceasy_challenge], HellaSwag [zellers2019hellaswag], PIQA [bisk2020piqa], and WinoGrande [sakaguchi2021winogrande].
Baselines. Since CryptoMoE is the first framework enabling private MoE inference, we compare it with three baselines: 1) Insecure baseline, 2) Dense baseline, 3) CryptoMoE baseline where $\Pi_{\text{dispatch}}$ and $\Pi_{\text{combine}}$ are constructed by CipherPrune [zhang2025cipherprune]’s pruning protocol.